# Meta Pruning via Graph Metanetworks : A Meta Learning Framework for Network Pruning

### v1

@@ 4.1 Setup
4.2 ResNet56 on CIFAR10
Table 1: ResNet56 on CIFAR10 (see full table in appendix 4)
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
FPGM [19] 93.59 92.93 -0.66 52.6% 2.11
FPC [21] 93.59 93.24 -0.35 52.9% 2.12
DMC [11] 93.62 92.69 -0.93 50.0% 2.00
GNN-RL [70] 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL [9] 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL [9] 93.53 93.77 0.24 52.6% 2.11
ATO [62] 93.50 93.74 0.24 55.0% 2.22
Meta-Pruning (ours) 93.51 94.08 0.57 56.5% 2.30
C-SGD [7] 93.39 93.44 0.05 60.8% 2.55
Greg-1 [60] 93.36 93.18 -0.18 60.8% 2.55
Greg-2 [60] 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL [9] 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL [9] 93.53 93.64 0.11 61.1% 2.57
ATO [62] 93.50 93.48 -0.02 65.3% 2.88
Meta-pruning (ours) 93.51 93.69 0.18 65.6% 2.91
Meta-pruning (ours) 93.51 93.25 -0.26 66.9% 3.02
Table 1 shows our results of pruning ResNet56 [16] on CIFAR10 [31]. Our algorithm outperforms all previous works. We can prune 65.6% FLOPs (i.e. 2.91x Speed Up) without any loss of accuracy, which is quite amazing. See Appendix B for more results and implementation details.
4.3 VGG19 on CIFAR100
Table 2: VGG19 on CIFAR100 (see full table in appendix 7)
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
OBD [59] 73.34 60.66 -12.68 83.58% 6.09
EigenD [59] 73.34 65.18 -8.16 88.64% 8.80
Greg-2 [60] 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL [9] 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL [9] 73.50 70.39 -3.11 88.79% 8.92
Meta-Pruning (ours) 73.65 70.06 -3.59 88.85% 8.97
Table 2 shows our results of pruning VGG19 [54] on CIFAR100 [31]. Our algorithm reaches 70.06% pruned accuracy with 88.85% FLOPs pruned (i.e. 8.97x Speed Up). While our pruning speed up is larger than all previous works, our pruned accuracy is only a little bit lower than DepGraph with SL and higher than all other works. See Appendix C for more results and implementation details.
4.4 ResNet50 on ImageNet
Table 3: ResNet50 on ImageNet (see full table in appendix 10)
DCP [77] 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP [49] 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM [19] 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP [42] 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC [11] 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random [38] 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph [9] 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO [62] 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP [39] 76.13%(-) 75.55%(-0.58%) - 56.7%
ours 76.14%(93.11%) 76.17%(+0.03%) 92.94%(-0.17%) 56.9%
Table 3 shows our results of pruning ResNet50 [16] on ImageNet[4]. Our algorithm reaches 76.17% top-1 and 92,95% top-2 accuracy with 56.9% FLOPs pruned. With the largest pruned FLOPs, our pruned accuracy is higher than almost all previous works (including DTP [39]). The only exception is that our accuracy is a bit lower than ATO [62], but our pruned FLOPs is much larger (1.7%) than it. See appendix D for more results and implementation details.
4.5 Ablation Study
How metanetwork works. To help readers visually grasp the effect of metanetwork, we visualized the "Test Acc VS. Speed Up" curve of both origin network and network after feedforward through metanetwork and finetuning . We generate the curve by prune the network little by little. After each pruning step, we evaluated the model’s performance on the test set (no finetuning in this step) and recorded a data point representing the current speed up and corresponding test accuracy. Connecting these points forms a curve that illustrates how accuracy changes as the speed up increases. The "Test Acc VS. Speed Up" curve of ResNet on CIFAR10 can be seen in figure 3(a). Similar curves of VGG19 on CIFAR100 and ResNet50 on ImageNet can be seen in appendix E.1.
As shown in the "Test Accuracy vs. Speed Up" curves, as the pruning speed up increases, the accuracy drops at a significantly slower rate after applying the metanetwork and finetuning. Take figure 3(a) as an example (this is exactly the metanetwork we use in 4.2). For the original network, the accuracy decreases rapidly and nearly collapses before reaching a speed up of 1.50x, suggesting that directly using group l2 norm criterion 3.3 to prune the origin network removes many important parts of the network (i.e. the origin network is hard to prune under this criterion). However, after metanetwork and finetuning, pruning to 2.0x speed up without finetuning causes almost no drop in accuracy, suggesting that using group l2 norm criterion on this network can efficiently distinguish the unimportant parts of this network (i.e. the network after metanetwork and finetuning is much easier to prune under this criterion). This supports our claim in 3 that metanetwork can transform a neural network that is difficult to prune into one that is much easier to prune.
(a) Origin VS After Metanetwork
(b) Impact Of Training Epochs On Metanetworks
Figure 3: Visualization of how metanetworks effect the networks (ResNet56 on CIFAR10)
Trade-off between accuracy and speed up. We observe an interesting phenomenon during meta-training: as the number of training epochs increases, the metanetwork’s ability to make the target network easier to prune becomes stronger. However, this comes at the cost of its ability to maintain the accuracy. As shown in figure 3(b), with more training epochs, the "Test Accuracy vs. Speed Up" curve shifts downward, and the flat portion of the curve becomes longer. This characteristic allows us to adaptively meet different pruning requirements. If a higher level of pruning is desired and a moderate drop in accuracy is acceptable, a metanetwork trained for more epochs would be preferable. Conversely, if preserving accuracy is more important, a metanetwork trained for fewer epochs would be a better choice.
Finetuning after Metanetwork. See appendix E.2 for our study on the impact of the number of finetuning epochs after metanetwork on performance.
5 Conclusion
In this paper we proposes a brand new meta-learning framework for network pruning. For the first time we introduce the concept of metanetwork (a network that takes another network as input and produces a modified network as output) into pruning. The metanetwork is trained to automatically learn effective pruning strategies. After training, it can take a network that is hard to prune as input and generate a transformed network that is much easier to prune. During pruning, our method needs nothing more than a feedforward through metanetwork and trivial finetuning to prune at state-of-the-art.
Advantages Over Previous Works. Our metanetwork is capable of considering the weights of multiple nearby layers, which sets it apart from most previous approaches. Unlike earlier methods that either focus only on local weights or incorporate more layers but rely on simple, hand-crafted relationships between them, our metanetwork automatically learns complex relationships among neighboring layers. These relationships are too intricate to be expressed by simple human-defined rules. So we use metanetwork to learn it automatically during training. This can explain why our work outperforms almost all previous works,
Limitations And Future Work. Our work also has several limitations. First, converting networks into graph representations requires GPUs with high VRAM, making it challenging to apply our method to large and densely parameterized models. Second, our current approach can’t be directly applied to networks with different architectures, such as Transformers or RNNs without appropriate modifications. However, these limitations also present opportunities for future research. We think the following directions are promising and worth exploring: (1) Extending our method to accommodate diverse architectures, including Transformers, diffusion models, and others. (2) Investigating alternative approaches to constructing the metanetworks. (3) Applying similar meta-learning frameworks to a broader range of tasks.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A General Experiment Setup
A.1 General Settings
Data for Meta-Training: As discussed in Section 3.4, meta-training requires two types of data: data models and traditional training datasets. To generate a data model, we train a base model, perform initial pruning followed by finetuning, and finally save the model along with relevant information. Typically, generating 3–10 data models is sufficient, with 2–8 used for meta-training and the remainder reserved for meta-evaluation or testing.
Visualizing Metanetworks: This is a key technique used throughout our experiments. When we refer to visualizing a metanetwork, we mean passing a model through the metanetwork, followed by finetuning, and then visualize the “Test Accuracy vs. Speed-Up” curve of the resulting model, as illustrated in Figure 3. This visualization provides insights into the metanetwork’s behavior and helps determine whether it is suitable for pruning. An ideal metanetwork should exhibit a long flat region in the curve where accuracy remains close to or above the target pruning accuracy. There are two ways to select a suitable metanetwork: (1) using meta-evaluation during meta-training, or (2) visualizing metanetworks post-training. In our final experiments, we exclusively use the second method, which is significantly more efficient. We typically do not visualize all metanetworks but instead search for the best one using a binary search strategy.
A.2 General Meta Training Details
Equivalent Conversion between Network and Graph: During meta-training, we first convert the original network into an origin graph. Then, we feed this origin graph through the metanetwork to generate a new graph. Finally, we convert the new graph back into a new network. When converting the original network into the origin graph, all network parameters are mapped to either node or edge features in the graph. However, some graph features do not correspond to any network parameter. For these missing features, we use default values that effectively simulate the absence of those parameters. When converting the new graph back into the new network, we only map those trainable parameters from the new graph back to the new network. For untrainable parameters in the new network, we keep it the same as the old network. We provide more explanations for this conversion process in the following subsections: B, C, and D.
Meta Training Milestone: As discussed in the Ablation Study 3, as the number of training epochs increases, the metanetwork’s ability to make the network easier to prune becomes stronger, while its ability to maintain the accuracy becomes weaker. During meta-training, we usually set a milestone for our learning rate scheduler. Before this milestone, we use a relatively large learning rate to quickly improve the metanetwork’s ability to make networks more prunable. After the milestone, we switch to a smaller learning rate to finetune the metanetwork and slow down the change of the two abilities, allowing us to select a well-balanced metanetwork. To determine a reasonable milestone value, we can first perform a short meta-training phase using the large learning rate, then visualize the resulting metanetworks to identify the point where the accuracy-maintaining ability of metanetwork is a bit stronger than we expect.
A.3 General Pruning Details :
Pruning Speed Up: Once a suitable metanetwork has been selected for pruning, the next step is to determine the target speed up. In general, after visualizing the metanetwork, we observe a flat region followed by a sharp decline in accuracy. We define a turning point as the speed up at which the accuracy drops below a certain threshold. Experimentally, we find that we can safely prune the model at a speed up slightly higher than this turning point without incurring any accuracy loss after fine-tuning. For instance, when pruning ResNet-56 on CIFAR-10, we consider an accuracy drop below 0.93 as the turning point (typically around 2.0$\times$ speed up), and in practice, we are able to achieve a 2.9$\times$ speed-up with no loss in accuracy after finetuning.
Appendix B ResNet56 on CIFAR10
B.1 Equivalent Conversion between Network and Graph
@@ B.2 Meta Training
Training: We train our metanetwork with learning rate 0.001, milestone "12", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 28 as our final network for pruning.
B.3 Full Results (Including Error Bars)
All results are summarized in Table 4, where we repeat the pruning process 3 times using different seeds: 7, 8, 9. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 1.3$\times$ speed up, the actual achieved speed up may be slightly larger, such as 1.31$\times$, 1.32$\times$, or 1.35$\times$.
@@ B.3 Full Results (Including Error Bars)
NISP [69] 43.2% 1.76
Geometric [20] 93.59 93.26 -0.33 41.2% 1.70
Polar [76] 93.80 93.83 0.03 46.8% 1.88
DCP-Adapt [77] 93.80 93.81 0.01 47.0% 1.89
CP [35] 92.80 91.80 -1.00 50.0% 2.00
AMC [23] 92.80 91.90 -0.90 50.0% 2.00
HRank [41] 93.26 92.17 -1.09 50.0% 2.00
SFP [18] 93.59 93.36 -0.23 52.6% 2.11
ResRep [8] 93.71 93.71 0.00 52.8% 2.12
SCP [28] 93.69 93.23 -0.46 51.5% 2.06
FPGM [19] 93.59 92.93 -0.66 52.6% 2.11
FPC [21] 93.59 93.24 -0.35 52.9% 2.12
DMC [11] 93.62 92.69 -0.93 50.0% 2.00
GNN-RL [70] 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL [9] 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL [9] 93.53 93.77 0.24 52.6% 2.11
ATO [62] 93.50 93.74 0.24 55.0% 2.22
Meta-Pruning (ours) 93.51 94.08 0.57 56.5% 2.30
Meta-Pruning (ours) 93.51 93.56 0.05 56.5% 2.30
Meta-Pruning (ours) 93.51 93.81 0.30 56.9% 2.32
GBN [67] 93.10 92.77 -0.33 60.2% 2.51
AFP [6] 93.93 92.94 -0.99 60.9% 2.56
C-SGD [7] 93.39 93.44 0.05 60.8% 2.55
Greg-1 [60] 93.36 93.18 -0.18 60.8% 2.55
Greg-2 [60] 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL [9] 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL [9] 93.53 93.64 0.11 61.1% 2.57
ATO [62] 93.50 93.48 -0.02 65.3% 2.88
Meta-pruning (ours) 93.51 93.69 0.18 65.6% 2.91
Meta-pruning (ours) 93.51 93.42 -0.09 66.0% 2.94
Meta-pruning (ours) 93.51 93.31 -0.20 66.6% 2.99
Meta-pruning (ours) 93.51 93.25 -0.26 66.9% 3.02
Meta-pruning (ours) 93.51 93.20 -0.31 66.7% 3.00
Meta-pruning (ours) 93.51 93.18 -0.33 67.5% 3.08
@@ B.4 Computer Resources and Hyperparameters
Final Pruning metanetwork epoch 28
@@
Appendix C VGG19 on CIFAR100
C.1 Equivalent Conversion between Network and Graph
@@ C.2 Meta Training
Training: We train our metanetwork with learning rate 0.001, milestone "15", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 51 as our final network for pruning.
C.3 Full Results (Including Error Bars)
All results are summarized in Table 7, where we repeat the pruning process 3 times using different seeds: 7, 8, 9. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 8.95$\times$ speed up, the actual achieved speed up may be slightly larger, such as 8.95$\times$, 9.01$\times$, or 9.02$\times$.
@@ C.3 Full Results (Including Error Bars)
OBD [59] 73.34 60.70 -12.64 82.55% 5.73
OBD [59] 73.34 60.66 -12.68 83.58% 6.09
EigenD [59] 73.34 65.18 -8.16 88.64% 8.80
Greg-1 [60] 74.02 67.55 -6.67 88.69% 8.84
Greg-2 [60] 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL [9] 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL [9] 73.50 70.39 -3.11 88.79% 8.92
Meta-Pruning (ours) 73.65 70.06 -3.59 88.85% 8.97
Meta-Pruning (ours) 73.65 69.19 -4.46 88.91% 9.02
Meta-Pruning (ours) 73.65 67.05 -6.60 89.07% 9.15
@@ C.4 Computer Resources and Hyperparameters
node res ratio 0.01
edge res ratio 0.01
@@
Final Pruning metanetwork epoch 51
speed up 8.95
@@
Appendix D ResNet50 on ImageNet
D.1 Equivalent Conversion between Network and Graph
@@
The edge features are constructed by simply flatten convolutional kernels into feature vectors (We have convolutional kernels of size $1\times 1$, $3\times 3$ or $7\times 7$). We treat all residual connections as convolutional layers with kernel size $1\times 1$. For example, consider two layers A and B connected via a residual connection, with neurons indexed as $1,2,3,\ldots$ in both layers. If the residual connection has no learnable parameters (i.e., it directly adds the input to the output), we represent the edge from $A_{i}$ to $B_{i}$ as a $1\times 1$ convolutional kernel with value 1. Different from section B, to reduce the VRAM footprint of the program, we build no edges between $A_{i}$ and $B_{j}$ ($i\neq j$). If the residual connection includes parameters (e.g., a downsample with a convolutional layer and batch normalization), we construct the edge features in the same way as for standard convolutional layers.
D.2 Meta Training
@@
One Meta Training Epoch : We train with torch.nn.parallel.DistributedDataParallel (pytorch data parallel) across 8 gpus. Unlike in Section B and Section C, at the beginning of each epoch, we evenly distribute different data models across the GPUs. For example, since we have two data models, we load one on four GPUs and the other on the remaining four GPUs. At each iteration, we forward all eight models (replicated across the 8 GPUs) through the metanetwork to generate eight new models. We then feedforward a large batch of ImageNet data, evenly distributed through these eight new models, to compute the accuracy loss. We also use the parameters of the eight new networks to compute the sparsity loss. Then we backward the gradients from both two losses to update our metanetwork. Each training epoch consists of (ImageNet data num / big batch size) iterations, meaning that one full pass over the entire ImageNet dataset when computing the accuracy loss is considered as one epoch .
Training: We train our metanetwork with learning rate 0.01, milestone "4", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 13 as our final network for pruning.
D.3 Full Results (Including Error Bars)
All results are summarized in Table 10, where we repeat the pruning process 3 times. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 2.3$\times$ speed up, the actual achieved speed up may be slightly larger, such as 2.31$\times$, 2.32$\times$, or 2.35$\times$.

### v2

@@ D.1.3 General Pruning Details :
D.2 ResNet56 on CIFAR10
D.2.1 Equivalent Conversion between Network and Graph
@@ D.2.2 Meta Training
Training: We train our metanetwork with learning rate 0.001, milestone "3", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 39 as our final network for pruning.
D.2.3 GPU Usage
All tasks can be run on 1 NVIDIA RTX 4090.
D.2.4 Full Results
[…]
@@ D.2.5 Hyperparameters
NISP (Yu et al., 2018) 43.2% 1.76
Geometric (He et al., 2019b) 93.59 93.26 -0.33 41.2% 1.70
Polar (Zhuang et al., 2020) 93.80 93.83 0.03 46.8% 1.88
DCP-Adapt (Zhuang et al., 2018) 93.80 93.81 0.01 47.0% 1.89
CP (Li et al., 2017a) 92.80 91.80 -1.00 50.0% 2.00
AMC (He et al., 2018b) 92.80 91.90 -0.90 50.0% 2.00
HRank (Lin et al., 2020a) 93.26 92.17 -1.09 50.0% 2.00
SFP (He et al., 2018a) 93.59 93.36 -0.23 52.6% 2.11
ResRep (Ding et al., 2021) 93.71 93.71 0.00 52.8% 2.12
SCP (Kang & Han, 2020) 93.69 93.23 -0.46 51.5% 2.06
FPGM (He et al., 2019a) 93.59 92.93 -0.66 52.6% 2.11
FPC (He et al., 2020) 93.59 93.24 -0.35 52.9% 2.12
DMC (Gao et al., 2020) 93.62 92.69 -0.93 50.0% 2.00
GNN-RL (Yu et al., 2022) 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL (Fang et al., 2023) 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL (Fang et al., 2023) 93.53 93.77 0.24 52.6% 2.11
ATO (Wu et al., 2024) 93.50 93.74 0.24 55.0% 2.22
Meta-Pruning (ours) 93.51 93.64 0.13 56.5% 2.30
Meta-Pruning (ours) 93.51 93.75 0.24 58.0% 2.38
Meta-Pruning (ours) 93.51 93.78 0.27 56.5% 2.30
GBN (You et al., 2019a) 93.10 92.77 -0.33 60.2% 2.51
AFP (Ding et al., 2018) 93.93 92.94 -0.99 60.9% 2.56
C-SGD (Ding et al., 2019) 93.39 93.44 0.05 60.8% 2.55
Greg-1 (Wang et al., 2021) 93.36 93.18 -0.18 60.8% 2.55
Greg-2 (Wang et al., 2021) 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL (Fang et al., 2023) 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL (Fang et al., 2023) 93.53 93.64 0.11 61.1% 2.57
ATO (Wu et al., 2024) 93.50 93.48 -0.02 65.3% 2.88
Meta-pruning (ours) 93.51 93.64 0.13 65.6% 2.91
Meta-pruning (ours) 93.51 93.28 -0.23 65.9% 2.93
Meta-pruning (ours) 93.51 93.49 -0.02 66.0% 2.94
Meta-pruning (ours) 93.51 93.27 -0.24 66.8% 3.01
Meta-pruning (ours) 93.51 93.10 -0.41 66.9% 3.02
Meta-pruning (ours) 93.51 93.52 0.01 67.0% 3.03
@@
Final Pruning metanetwork epoch 39
@@
D.3 VGG19 on CIFAR100
D.3.1 Equivalent Conversion between Network and Graph
@@ D.3.2 Meta Training
Training: We train our metanetwork with learning rate 0.001, milestone "10", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 38 as our final network for pruning.
D.3.3 GPU Usage
All tasks can be run on 1 NVIDIA RTX 4090.
D.3.4 Full Results
[…]
@@ D.3.5 Hyperparameters
OBD (Wang et al., 2019) 73.34 60.70 -12.64 82.55% 5.73
OBD (Wang et al., 2019) 73.34 60.66 -12.68 83.58% 6.09
EigenD (Wang et al., 2019) 73.34 65.18 -8.16 88.64% 8.80
Greg-1 (Wang et al., 2021) 74.02 67.55 -6.67 88.69% 8.84
Greg-2 (Wang et al., 2021) 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL (Fang et al., 2023) 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL (Fang et al., 2023) 73.50 70.39 -3.11 88.79% 8.92
Meta-Pruning (ours) 73.65 68.65 -5.00 88.81% 8.94
Meta-Pruning (ours) 73.65 67.63 -6.02 88.96% 9.06
Meta-Pruning (ours) 73.65 69.75 -3.90 88.83% 8.95
@@
node res ratio 0.05
edge res ratio 0.05
@@
Final Pruning metanetwork epoch 38
speed up 8.90
@@
D.4 ResNet50 on ImageNet
D.4.1 Equivalent Conversion between Network and Graph
@@
The edge features are constructed by simply flatten convolutional kernels into feature vectors (We have convolutional kernels of size $1\times 1$, $3\times 3$ or $7\times 7$). We treat all residual connections as convolutional layers with kernel size $1\times 1$. For example, consider two layers A and B connected via a residual connection, with neurons indexed as $1,2,3,\ldots$ in both layers. If the residual connection has no learnable parameters (i.e., it directly adds the input to the output), we represent the edge from $A_{i}$ to $B_{i}$ as a $1\times 1$ convolutional kernel with value 1. Different from section D.2, to reduce the VRAM footprint of the program, we build no edges between $A_{i}$ and $B_{j}$ ($i\neq j$). If the residual connection includes parameters (e.g., a downsample with a convolutional layer and batch normalization), we construct the edge features in the same way as for standard convolutional layers.
D.4.2 Meta Training
@@
One Meta Training Epoch : We train with torch.nn.parallel.DistributedDataParallel (pytorch data parallel) across 8 gpus. Unlike in Section D.2 and Section D.3, at the beginning of each epoch, we evenly distribute different data models across the GPUs. For example, since we have two data models, we load one on four GPUs and the other on the remaining four GPUs. At each iteration, we forward all eight models (replicated across the 8 GPUs) through the metanetwork to generate eight new models. We then feedforward a large batch of ImageNet data, evenly distributed through these eight new models, to compute the accuracy loss. We also use the parameters of the eight new networks to compute the sparsity loss. Then we backward the gradients from both two losses to update our metanetwork. Each training epoch consists of (ImageNet data num / big batch size) iterations, meaning that one full pass over the entire ImageNet dataset when computing the accuracy loss is considered as one epoch .
Training: We train our metanetwork with learning rate 0.01, milestone "2", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 12 as our final network for pruning.
D.4.3 GPU Usage
Meta-Training and feed forward through the metanetwork during pruning requires large VRAM and needs to be run on NVIDIA A100. All other training and finetuning can be run on NVIDIA RTX 4090. In practice, we use 8 gpus in parallel.
D.4.4 Full Results
[…]